# Learning Group Activity Features Through Person Attribute Prediction

### 05 Mar 2024 v1

@@ 1 Introduction
Group Activity Recognition (GAR), in which a frame or video is classified into either of the predefined group activity classes, has been widely investigated [14, 36, 2, 24, 6, 9, 40, 19, 11, 28, 43, 16, 3, 37]. All of these GAR methods are based on supervised learning that requires the ground-truth group activities, as shown in Fig. 1 (a). In addition, it is known that person action recognition, which is also achieved based on a supervised learning manner, supports GAR [14, 36, 2, 24, 6, 9, 19, 11, 28, 43, 3, 38].
@@ 2.1 Group Activity Recognition
Supervision by the labels of person action and group activity. Different from the aforementioned methods only with group activity supervision, the GAR network is jointly trained with the person action recognition network in [14, 36, 2, 24, 6, 9, 19, 11, 28, 43, 3] for augmenting GAR. In [36, 6], GNN models the interactions between person features. In [9, 19, 11, 28, 43], Transformer [30] improves modeling the interactions between person features for GAR.

### 11 Mar 2024 v2

@@ 1 Introduction
Group Activity Recognition (GAR), in which a frame or video is classified into either of the predefined group activity classes, has been widely investigated in recent years  [14, 37, 2, 25, 6, 9, 42, 19, 11, 29, 45, 16, 3, 39, 21, 38, 40]. All of these GAR methods are based on supervised learning that requires the ground-truth group activities, as shown in Fig. 1 (a). In addition, it is known that person action recognition, which is also achieved based on a supervised learning manner, supports GAR [14, 37, 2, 25, 6, 9, 19, 11, 29, 45, 3, 40, 21, 38].
@@ 2.1 Group Activity Recognition
Supervision by the labels of person action and group activity. Different from the aforementioned methods only with group activity supervision, the GAR network is jointly trained with the person action recognition network in [14, 37, 2, 25, 6, 9, 19, 11, 29, 45, 3, 21] for augmenting GAR. In [37, 6], GNN models the interactions between person features. In [9, 19, 11, 29, 45], Transformer [31] improves modeling the interactions between person features for GAR.